OpenStudy (anonymous):

S.western Moving and Storage wants to have enough money to purchase a new tractor-trailor in 5 years at a cost of $290,000. If the company sets aside $100,000 in year 2 and $75,000 in year 3, how much will the company need to set aside in year 4 in order to have the money it needs if the money set aside earns 9% per year?

OpenStudy (anonymous):

I got an answer of 65495.05

OpenStudy (anonymous):

It is possible I may have made slight rounding errors but the basic equation you will set-up is simple algebra as follows: \[290,000=x+(x*.09)+[75,000+75,000*.09]*(1+.09) + {100,000+100,000*.09} *(1+.09)^{2}\]

OpenStudy (anonymous):

for example using single payments w/ 9% - compound amount n=3... (f/p)= 1.2950 and n=2...(f/p)= 1.1881 and present worth n=1...(p/f)= .9174)

OpenStudy (anonymous):

I would think you could just use a simple compound interest formula actually. That being set up as follow since FV= PV(1+i)^n where i is interest and n is duration \[290,000= x(1+.09)^{1}+75,000(1+.09)^{2}+100,000(1+.09)^3\]